OpenStudy (anonymous):

Factor the polyomial completely using any method 2z^4-1250 Medal given for answer

OpenStudy (ybarrap):

2(x^4 - 625) 2( (x^2)^2 - 25^2) 2( x^2 - 25)(x^2 + 25) 2(x+5)(X-5)(x^2 + 25)

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

the final ans is 2(x+5)(X-5)(x^2 + 25) as given above by ybarrap

OpenStudy (anonymous):

Thanks can you help me with this on too ? 6b^4-17b^2-28

OpenStudy (anonymous):

pls check if u hv typed it correctly.......

OpenStudy (anonymous):

nope its right

OpenStudy (ybarrap):

let x = b^2 6b^4-17b^2-28 =6x^2 - 17x - 28 (6x + 7)(6x -24 )/6 =(6x + 7)6(x - 4)/6 =(6x + 7)(x -4) substituting back: (6b^2 + 7)(b^2 -4) (6b^2 + 7)(b + 2)(b - 2) Ans. (6b^2 + 7)(b+2)(b-2)